OpenStudy (anonymous):

solve (5x+4)^1/2-3x=0

OpenStudy (anonymous):

Draw it out. It's better

OpenStudy (anonymous):

\[(5x+4)^\frac{1}{2}-3x=0\] \[\sqrt{(5x+4)}=3x\] Squaring both sides we find \[5x+4= 9x^2 \rightarrow 9x^2-5x -4 =0\] \[9x^2-9x +4x -4 =0 \rightarrow 9x(x-1) +4(x -1) =0\] (9x+4)(x -1) =0 i.e. (9x+4) =0 or (x -1)=0 x = 1 , -4/9 => reject x = -4/9 since it won't work in the original equation x = 1 => answer